# pulsars; FRBs

### 08 Apr 2024 v1

@@ Abstract
We present the first new type of solution of the pulsar equation since 1999. In it, the whole magnetosphere is confined inside the light cylinder and an electrically charged layer wraps around it and holds it together. The reason this new solution has never been obtained before is that all current time-dependent simulations are initialized with a vacuum dipole configuration that extends to infinity, thus their final steady-state solution also extends to infinity. Under special conditions, such a confined configuration may be attained when the neutron star first forms in the interior of a collapsing star during a supernova explosion, or when it accretes from an external wind or disk from a donor star. It is shown that this new maximally closed non-decelerating solution is the limit of a continuous sequence of standard magnetospheres with open and closed field lines when the amount of open field lines gradually drops to zero. The minimum energy solution in this sequence is a standard magnetosphere in which the closed field line region extends up to about $80\%$ of the light cylinder. We estimate that the released energy when the new solution transitions to the minimum energy one is enough to power a fast radio burst.
@@ 1 The pulsar equation
Right after the discovery of pulsars in 1968 Hewish et al. (1969), Goldreich and Julian were the first to sketch the general structure of the pulsar magnetosphere Goldreich and Julian (1969) . The detailed structure of the axisymmetric pulsar magnetosphere was obtained much later by Contopoulos, Kazanas and Fendt who clearly identified the presence of a global magnetospheric poloidal electric current sheet Contopoulos et al. (1999). Since then, several authors have addressed the pulsar magnetosphere problem with Force-Free Electrodynamic (FFE) simulations (Spitkovsky, 2006; Kalapotharakos and Contopoulos, 2009, etc.), Magnetohydrodynamic (MHD) simulations (Tchekhovskoy et al., 2013, etc.), ‘ab initio’ Particle-In-Cell (PIC) simulations (Philippov and Spitkovsky, 2014; Philippov et al., 2015; Kalapotharakos et al., 2018, etc.), and most recently with Machine Learning Stefanou et al. (2023); Dimitropoulos et al. (2024). Several inconsistencies in recent state-of-the-art numerical simulations (e.g. the extent of the co-rotating closed-line region, the microscopic treatment and the thickness of the current sheet, magnetospheric dissipation, etc.) led us to believe that one can only trust them qualitatively (not quantitatively) to make meaningful comparisons with observations. This is why, in Dimitropoulos et al. (2024) we proposed to return to the basics and obtain the reference ideal steady-state force-free magnetosphere with a novel independent way, namely with Machine Learning.
@@
where ${\bf E}$ and $\bf{B}$ are the electric and magnetic fields respectively, $\rho_{e}=\nabla\cdot{\bf E}/(4\pi)$ is the electric charge density, and ${\bf J}$ is the electric current density. Under axisymmetric conditions one may define the magnetic flux function $\Psi$ such that $B_{r}=\partial\Psi/\partial\theta/(r^{2}\sin\theta)$ and $B_{\theta}=-\partial\Psi/\partial r/(r\sin\theta)$ which automatically satisfies the condition $\nabla\cdot{\bf B}=0$. Under steady-state axisymmetric force-free ideal conditions, eq. (1) becomes the well known pulsar equation, namely
$\left(1-\frac{r^{2}\sin^{2}\theta}{R_{\rm LC}^{2}}\right)\left[\frac{\partial^{2}\Psi}{\partial r^{2}}-\frac{\partial\Psi}{\partial\theta}\frac{\cos\theta}{r^{2}\sin\theta}+\frac{1}{r^{2}}\frac{\partial^{2}\Psi}{\partial\theta^{2}}\right]-\frac{2r\sin\theta}{R_{\rm LC}^{2}}\left[\frac{\partial\Psi}{\partial\theta}\frac{\cos\theta}{r}+\frac{\partial\Psi}{\partial r}\sin\theta\right]$
$+II^{\prime}(\Psi)=0$ (2)
in spherical coordinates $r,\theta$ Scharlemann and Wagoner (1973). Here, $I=I(\Psi)$ is the distribution of poloidal electric current from which one obtains the azimuthal component of the magnetic field $B_{\phi}=I/(r\sin\theta)$, and $I^{\prime}\equiv{\rm d}I/{\rm d}\Psi$. Eq. (2) is an elliptical partial differential equation with a singularity along the light cylinder, the nominal distance where the co-rotational velocity becomes equal to the speed of light ($R_{\rm LC}\equiv c/\Omega$, where $\Omega$ is the angular velocity of stellar rotation). This equation was first solved with a special numerical technique developed by Contopoulos et al. (1999) that was used again since then by several others (e.g. Gruzinov (2005); Timokhin (2006); Ntotsikas et al. (2024)).
In § 2 we present a new type of solution of eq. (2) that is very different from the canonical solutions found in the literature. In § 3 we obtain new solutions of the ideal force-free pulsar magnetosphere for various amounts of open magnetic flux using the novel Machine Learning methodology of Dimitropoulos et al. (2024) (hereafter Paper I; see description in Appendix 1). We show that the new type of solution is the limit of a continuous sequence of standard magnetospheres with open and closed field lines in which the closed line region approaches closer and closer to the light cylionder and the amount of open field lines gradually drops to zero. In § 4 we discuss the possibility that an abrupt transition from our new magnetospheric solution to a standard magnetosphere with open and closed field lines may generate a Fast Radio Burst. We conclude in § 5 with a discussion of our results. We consider only axisymmetric magnetospheres, and in a future work we will investigate whether our results apply also to an oblique rotator.
@@ 2 A confined magnetosphere
Let us consider here a theoretical limit that has never been considered before, namely one with no open field lines and no poloidal electric current ($I=0$). In that limit, the whole magnetosphere of the central dipole is contained within the light cylinder, with boundary conditions $\Psi(r,\theta=0)=\Psi(r,\theta=\pi)=0$, a dipole field $\Psi(r_{*},\theta)=\Psi_{\rm max}\sin^{2}\theta$ along the surface of the central star ($\Psi_{\rm max}$ is the total dipolar magnetic flux that emanates from each stellar hemisphere), and $\Psi(r=R_{\rm LC}/\sin\theta,\theta)=0$ along the light cylinder. The solution of eq. (2) with these boundary conditions is obtained with the method described in Gourgouliatos and Lynden-Bell (2008) and is shown in figure 1. This configuration co-rotates with the central star and is wrapped around by an electrically charged layer at the light cylinder. It is a unique feature of special relativity that allows such a spatially confined electromagnetic field solution, namely the presence of electric charges and electric fields that are generated by the rigid co-rotation of the magnetosphere. In the region next to the light cylinder in particular, the electric charge density is positive/negative and generates (via its co-rotation) an azimuthal electric current along/opposite to the direction of rotation for an aligned/counter-aligned rotator respectively. The magnetic field points downwards/upwards and the electric field outwards/inwards respectively, so it is obvious that the inward Lorentz force ${\bf J}\times{\bf B}/c$ is balanced by the outward electrostatic force $\rho_{e}{\bf E}$.
It is, however, interesting that such a confined solution with vacuum outside is not possible without a light cylinder, as is the case in non-relativistic MHD. In particular, pressure balance across the electrically charged current sheet along the light cylinder may be written as
$\left.(B^{2}-E^{2})\right|_{\rm IN}=\left.(B^{2}-E^{2})\right|_{\rm OUT}=0\ ,$ (3)
@@
$\left.B_{z}^{2}(1-x^{2})\right|_{\rm IN}=0\ .$ (4)
The requirement for the continuity of $B^{2}-E^{2}$ across an infinitely thin relativistic current sheet, eq. (3), stems from the integration of the force-free equation $\rho_{e}{\bf E}+{\bf J}\times{\bf B}/c={\bf 0}$ across it Lyubarsky (2005); Uzdensky (2003). Here, $x\equiv R/R_{\rm LC}$ is the cylindrical radius in units of the light cylinder radius $R_{\rm LC}$, and $\left.x\right|_{\rm IN}\rightarrow 1^{-}$. It is assumed that the whole magnetosphere is confined inside the light cylinder, it remains untwisted (i.e. $B_{\phi}=0$), it co-rotates with the central star, and there is nothing outside. Obviously, in the limit $\left.x\right|_{\rm IN}=1^{-}$, pressure balance (i.e. eq. 4) is satisfied, thus solutions spatially confined inside the light cylinder are possible. In this new solution, the value of $\left.B_{z}\right|_{\rm IN}$ is non-zero and can be determined numerically. We emphasize once again that such spatially self-confined solutions with vacuum outside are possible only along the light cylinder, not inside, nor outside.
@@
We propose that this singular configuration may be realized in nature under special conditions during the stellar collapse that led to the formation of a spinning neutron star in a supernova explosion. The initial neutron star magnetosphere will be squeezed to a thin layer above the surface of the neutron star and will rotate with the neutron star. The solutions of the pulsar equation shown in figure 2 were obtained with the same method by setting $\Psi=0$ along a sphere of radius $r_{\rm ejecta}$, and along the light cylinder wherever $r_{\rm ejecta}\sin\theta>R_{\rm LC}$. Thus, the whole magnetosphere is confined between the neutron star surface and a spherical shell of supernova ejecta. As the supernova explosion proceeds and the ejecta expand, the size $r_{\rm ejecta}$ of the magnetosphere will increase. During that stage, the whole magnetosphere co-rotates with the central neutron star, thus it will be confined inside the light cylinder wherever $r_{\rm ejecta}\sin\theta>R_{\rm LC}$. It is implied here that, during the supernova explosion, as $r_{\rm ejecta}$ grows slowly at subluminal velocities, the magnetosphere evolves fast into a sequence of steady-state solutions. When $r_{\rm ejecta}$ grows beyond $r_{\rm ejecta}=R_{\rm LC}$, any part of the magnetosphere that crosses the light cylinder cannot open up to infinity as in the magnetosphere of an isolated pulsar. Therefore, it will be wound very fast by the stellar rotation, the Lorentz force will squeeze it vertically, and it will eventually detach from the rest of the magnetosphere that remains confined inside the light cylinder as in the solutions shown in figure 2. It is obvious that this is an idealized situation in which a neutron star and its magnetosphere formed and remained at the center of the collapse (in general, the crust of a young neutron star may not have formed/solidified yet, and the neutron star would also potentially experience a kick that will move it away from the center of the collapse). It is also conceivable that such a confined configuration is attained when a neutron star accretes from a surrounding disk or from external winds from a donor star (e.g. Shakura and Sunyaev (1973); Doroshenko et al. (2011)). It is also obvious that the solutions shown in figures 1 and 2 are force-free approximations since they were obtained without taking into consideration the baryon rich and dense environment of a new born neutron star.
@@
The idealized solution shown in figure 1 is most probably unstable. We haven’t performed a formal stability analysis of the light cylinder boundary, but we understand this instability as follows: Whenever a part of the confined magnetosphere extends beyond the light cylinder, it cannot keep co-rotating with the central star without an azimuthal component of the magnetic field. In other words, when an initially poloidal magnetic field line crosses the light cylinder it is swept back so that the particle drift velocity
$v_{\rm D}$$\equiv$ $\frac{|{\bf E}\times{\bf B}|}{B^{2}}c=\frac{|{\bf E}_{p}\times{\bf B}_{p}+{\bf E}_{p}\times B_{\phi}\hat{\phi}|}{B_{p}^{2}+B_{\phi}^{2}}c$ (5)
$=$ $\frac{\sqrt{B_{p}^{4}x^{2}+B_{p}^{2}B_{\phi}^{2}x^{2}}}{B_{p}^{2}+B_{\phi}^{2}}c=\frac{B_{p}x}{\sqrt{B_{p}^{2}+B_{\phi}^{2}}}c<c$
@@
always remains subluminal. Without the azimuthal component $B_{\phi}$, $v_{D}>c$ outside the light cylinder where $x>1$. Therefore, any part of the confined magnetosphere that crosses the light cylinder will develop an azimuthal magnetic field component that will push it outwards towards infinity. Thus, the magnetosphere will transition to a standard solution with open and closed field lines, emitting an electromagnetic pulse of azimuthal magnetic field along the way. A similar effect is seen in numerical simulations during the evolution of initially poloidal magnetic field configurations when the stellar rotation is initiated (e.g. Spitkovsky, 2006). We will return to a further discussion of this burst in section 3 below. Notice, that using eq. (5) with $B_{\phi}=0$, the drift velocity is exactly equal to $c$ at the light cylinder, and the neglect of inertial terms is more questionable for the confined solution than for the standard one near the light cylinder. Inertial effects are even more pronounced near the light cylinder because, due to the $(1-x^{2})$ term in the denominator of the Goldreich-Julian charge density $\rho_{e}=-B_{z}/[2\pi R_{\rm LC}(1-x^{2})]$ Timokhin (2006), the volume integral of $\rho_{e}$ and the corresponding mass content of the idealized confined magnetosphere diverge. This suggests that, if realized in nature, the confined solution will be limited to within some short distance from the light cylinder.

### 11 Apr 2024 v2

@@ Abstract
We present the first new type of solution of the pulsar equation since 1999. In it, the whole magnetosphere is confined inside the light cylinder and an electrically charged layer wraps around it and holds it together. The reason this new solution has never been obtained before is that all current time-dependent simulations are initialized with a vacuum dipole configuration that extends to infinity; thus, their final steady-state solution also extends to infinity. Under special conditions, such a confined configuration may be attained when the neutron star first forms in the interior of a collapsing star during a supernova explosion, or when it accretes from an external wind or disk from a donor star. It is shown that this new maximally closed non-decelerating solution is the limit of a continuous sequence of standard magnetospheres with open and closed field lines when the amount of open field lines gradually drops to zero. The minimum energy solution in this sequence is a standard magnetosphere in which the closed field line region extends up to about $80\%$ of the light cylinder. We estimate that the released energy when the new solution transitions to the minimum energy one is enough to power a fast radio burst.
@@ 1 The Pulsar Equation
Right after the discovery of pulsars in 1968 Hewish et al. (1969), Goldreich and Julian were the first to sketch the general structure of the pulsar magnetosphere Goldreich and Julian (1969). The detailed structure of the axisymmetric pulsar magnetosphere was obtained much later by Contopoulos, Kazanas and Fendt, who clearly identified the presence of a global magnetospheric poloidal electric current sheet Contopoulos et al. (1999). Since then, several authors have addressed the pulsar magnetosphere problem with force-free electrodynamic (FFE) simulations (Spitkovsky (2006); Kalapotharakos and Contopoulos (2009), etc.) magnetohydrodynamic (MHD) simulations (Tchekhovskoy et al. (2013), etc.) ‘ab initio’ particle-in-cell (PIC) simulations (Philippov and Spitkovsky (2014); Philippov et al. (2015); Kalapotharakos et al. (2018), etc.) and most recently with Machine Learning Stefanou et al. (2023); Dimitropoulos et al. (2024). Several inconsistencies in recent state-of-the-art numerical simulations (e.g., the extent of the co-rotating closed-line region, the microscopic treatment and the thickness of the current sheet, magnetospheric dissipation, etc.) led us to believe that one can only trust them qualitatively (not quantitatively) to make meaningful comparisons with observations. This is why, in Dimitropoulos et al. (2024), we proposed to return to the basics and obtain the reference ideal steady-state force-free magnetosphere in a novel independent way, namely with Machine Learning.
@@
where ${\bf E}$ and $\bf{B}$ are the electric and magnetic fields, respectively, $\rho_{e}=\nabla\cdot{\bf E}/(4\pi)$ is the electric charge density, and ${\bf J}$ is the electric current density. Under axisymmetric conditions, one may define the magnetic flux function $\Psi$ such that ${\rm B}_{r}$ $=\partial\Psi/\partial\theta/(r^{2}\sin\theta)$ and ${\rm B}_{\theta}=-\partial\Psi/\partial r/(r\sin\theta)$, which automatically satisfies the condition $\nabla\cdot{\bf B}=0$. Under steady-state axisymmetric force-free ideal conditions, Equation (1) becomes the well-known pulsar equation, namely
$\left(1-\frac{r^{2}\sin^{2}\theta}{R_{\rm LC}^{2}}\right)\left[\frac{\partial^{2}\Psi}{\partial r^{2}}-\frac{\partial\Psi}{\partial\theta}\frac{\cos\theta}{r^{2}\sin\theta}+\frac{1}{r^{2}}\frac{\partial^{2}\Psi}{\partial\theta^{2}}\right]-\frac{2r\sin\theta}{R_{\rm LC}^{2}}\left[\frac{\partial\Psi}{\partial\theta}\frac{\cos\theta}{r}+\frac{\partial\Psi}{\partial r}\sin\theta\right]+II^{\prime}(\Psi)=0$ (2)
in spherical coordinates $r,\theta$ Scharlemann and Wagoner (1973). Here, $I=I(\Psi)$ is the distribution of poloidal electric current from which one obtains the azimuthal component of the magnetic field ${\rm B}_{\phi}=I/(r\sin\theta)$, and $I^{\prime}\equiv{\rm d}I/{\rm d}\Psi$. Equation (2) is an elliptical partial differential equation with a singularity along the light cylinder, the nominal distance where the co-rotational velocity becomes equal to the speed of light ($R_{\rm LC}\equiv c/\Omega$, where $\Omega$ is the angular velocity of stellar rotation). This equation was first solved with a special numerical technique developed by Contopoulos et al. (1999) that was used again since then by several others (e.g., Gruzinov (2005); Timokhin (2006); Ntotsikas et al. (2024)).
In Section 2, we present a new type of solution of Equation (2) that is very different from the canonical solutions found in the literature. In Section 3, we obtain new solutions of the ideal force-free pulsar magnetosphere for various amounts of open magnetic flux using the novel Machine Learning methodology of Dimitropoulos et al. (2024) (hereafter Paper I; see description in Appendix A). We show that the new type of solution is the limit of a continuous sequence of standard magnetospheres with open and closed field lines in which the closed line region approaches closer and closer to the light cylinder and the amount of open field lines gradually drops to zero. In Section 4 we discuss the possibility that an abrupt transition from our new magnetospheric solution to a standard magnetosphere with open and closed field lines may generate a fast radio burst. We conclude in Section 5 with a discussion of our results. We consider only axisymmetric magnetospheres, and in future work, we will investigate whether our results apply also to an oblique rotator.
@@ 2 A Confined Magnetosphere
Let us consider here a theoretical limit that has never been considered before, namely one with no open field lines and no poloidal electric current ($I=0$). In that limit, the whole magnetosphere of the central dipole is contained within the light cylinder, with boundary conditions $\Psi(r,\theta=0)=\Psi(r,\theta=\pi)=0$, a dipole field $\Psi(r_{*},\theta)=\Psi_{\rm max}\sin^{2}\theta$ along the surface of the central star ($\Psi_{\rm max}$ is the total dipolar magnetic flux that emanates from each stellar hemisphere), and $\Psi(r=R_{\rm LC}/\sin\theta,\theta)=0$ along the light cylinder. The solution of Equation (2) with these boundary conditions is obtained with the method described in Gourgouliatos and Lynden-Bell (2008) and is shown in Figure 1. This configuration co-rotates with the central star and is wrapped around by an electrically charged layer at the light cylinder. It is a unique feature of special relativity that allows such a spatially confined electromagnetic field solution, namely the presence of electric charges and electric fields that are generated by the rigid co-rotation of the magnetosphere. In the region next to the light cylinder in particular, the electric charge density is positive/negative and generates (via its co-rotation) an azimuthal electric current along/opposite to the direction of rotation for an aligned/counter-aligned rotator, respectively. The magnetic field points downwards/upwards and the electric field outwards/inwards, respectively, so it is obvious that the inward Lorentz force ${\bf J}\times{\bf B}/c$ is balanced by the outward electrostatic force $\rho_{e}{\bf E}$.
It is, however, interesting that such a confined solution with a vacuum outside is not possible without a light cylinder, as is the case in non-relativistic MHD. In particular, pressure balance across the electrically charged current sheet along the light cylinder may be written as
$\left.({\rm B}^{2}-{\rm E}^{2})\right|_{\rm IN}=\left.({\rm B}^{2}-{\rm E}^{2})\right|_{\rm OUT}=0\ ,$ (3)
@@
$\left.{\rm B}^{2}_{z}(1-x^{2})\right|_{\rm IN}=0\ .$ (4)
The requirement for the continuity of ${\rm B}^{2}-{\rm E}^{2}$ across an infinitely thin relativistic current sheet, Equation (3), stems from the integration of the force-free equation $\rho_{e}{\bf E}+{\bf J}\times{\bf B}/c={\bf 0}$ across it Lyubarsky (2005); Uzdensky (2003). Here, $x\equiv R/R_{\rm LC}$ is the cylindrical radius in units of the light cylinder radius $R_{\rm LC}$, and $\left.x\right|_{\rm IN}\rightarrow 1^{-}$. It is assumed that the whole magnetosphere is confined inside the light cylinder, it remains untwisted (i.e., ${\rm B}_{\phi}=0$), it co-rotates with the central star, and there is nothing outside. Obviously, in the limit $\left.x\right|_{\rm IN}=1^{-}$, pressure balance (i.e., Equation (4)) is satisfied; thus, solutions spatially confined inside the light cylinder are possible. In this new solution, the value of $\left.{\rm B}_{z}\right|_{\rm IN}$ is non-zero and can be determined numerically. We emphasize once again that such spatially self-confined solutions with a vacuum outside are possible only along the light cylinder, not inside or outside.
@@
We propose that this singular configuration may be realized in nature under special conditions during the stellar collapse that led to the formation of a spinning neutron star in a supernova explosion. The initial neutron star magnetosphere will be squeezed to a thin layer above the surface of the neutron star and will rotate with the neutron star. The solutions of the pulsar equation shown in Figure 2 were obtained with the same method by setting $\Psi=0$ along a sphere of radius $r_{\rm ejecta}$, and along the light cylinder wherever $r_{\rm ejecta}\sin\theta>R_{\rm LC}$. Thus, the whole magnetosphere is confined between the neutron star surface and a spherical shell of supernova ejecta. As the supernova explosion proceeds and the ejecta expand, the size $r_{\rm ejecta}$ of the magnetosphere will increase. During that stage, the whole magnetosphere co-rotates with the central neutron star; thus, it will be confined inside the light cylinder wherever $r_{\rm ejecta}\sin\theta>R_{\rm LC}$. It is implied here that, during the supernova explosion, as $r_{\rm ejecta}$ grows slowly at subluminal velocities, the magnetosphere evolves fast into a sequence of steady-state solutions. When $r_{\rm ejecta}$ grows beyond $r_{\rm ejecta}=R_{\rm LC}$, any part of the magnetosphere that crosses the light cylinder cannot open up to infinity as in the magnetosphere of an isolated pulsar. Therefore, it will be wound very fast by the stellar rotation, the Lorentz force will squeeze it vertically, and it will eventually detach from the rest of the magnetosphere that remains confined inside the light cylinder as in the solutions shown in Figure 2. It is obvious that this is an idealized situation in which a neutron star and its magnetosphere formed and remained at the center of the collapse (in general, the crust of a young neutron star may not have formed/solidified yet, and the neutron star would also potentially experience a kick that will move it away from the center of the collapse). It is also conceivable that such a confined configuration is attained when a neutron star accretes from a surrounding disk or from external winds from a donor star (e.g., Shakura and Sunyaev (1973); Doroshenko et al. (2011)). It is also obvious that the solutions shown in Figures 1 and 2 are force-free approximations since they were obtained without taking into consideration the baryon-rich and dense environment of a newborn neutron star.
@@
The idealized solution shown in Figure 1 is most probably unstable. We have not performed a formal stability analysis of the light cylinder boundary, but we understand this instability as follows: Whenever a part of the confined magnetosphere extends beyond the light cylinder, it cannot keep co-rotating with the central star without an azimuthal component of the magnetic field. In other words, when an initially poloidal magnetic field line crosses the light cylinder, it is swept back so that the particle drift velocity
$v_{\rm D}$$\equiv$ $\frac{|{\bf E}\times{\bf B}|}{{\rm B}^{2}}c=\frac{|{\bf E}_{p}\times{\bf B}_{p}+{\bf E}_{p}\times{\rm B}_{\phi}\hat{\phi}|}{{\rm B}_{p}^{2}+{\rm B}_{\phi}^{2}}c$ (5)
$=$ $\frac{\sqrt{{\rm B}_{p}^{4}x^{2}+{\rm B}_{p}^{2}{\rm B}_{\phi}^{2}x^{2}}}{{\rm B}_{p}^{2}+{\rm B}_{\phi}^{2}}c=\frac{{\rm B}_{p}x}{\sqrt{{\rm B}_{p}^{2}+{\rm B}_{\phi}^{2}}}c<c$
@@
always remains subluminal. Without the azimuthal component ${\rm B}_{\phi}$, $v_{D}>c$ outside the light cylinder, where $x>1$. Therefore, any part of the confined magnetosphere that crosses the light cylinder will develop an azimuthal magnetic field component that will push it outwards toward infinity. Thus, the magnetosphere will transition to a standard solution with open and closed field lines, emitting an electromagnetic pulse of azimuthal magnetic field along the way. A similar effect is seen in numerical simulations during the evolution of initially poloidal magnetic field configurations when the stellar rotation is initiated (e.g., Spitkovsky (2006)). We will return to a further discussion of this burst in Section 3 below. Notice that using Equation (5) with ${\rm B}_{\phi}=0$, the drift velocity is exactly equal to $c$ at the light cylinder, and the neglect of inertial terms is more questionable for the confined solution than for the standard one near the light cylinder. Inertial effects are even more pronounced near the light cylinder because, due to the $(1-x^{2})$ term in the denominator of the Goldreich–Julian charge density $\rho_{e}=-{\rm B}_{z}/[2\pi R_{\rm LC}(1-x^{2})]$ Timokhin (2006), the volume integral of $\rho_{e}$ and the corresponding mass content of the idealized confined magnetosphere diverge. This suggests that, if realized in nature, the confined solution will be limited to within some short distance from the light cylinder.